# Sayantan Roy Sankha S. Basu Mihir K. Chakraborty School of Cognitive Science

### v1

@@ 3. $\bullet$
    Let $(\mathscr{L},W)$ be a logical structure induced by a non-reflexive $q$-consequence operator. Note that, by Theorem 3.8, it is functionally $3$-valued. Suppose that it is not inferentially $3$-valued. So, there exists some functionally $2$-valued semantics for $\mathscr{L}$, say $\mathfrak{S}=(\textbf{M},\models_{1},\models_{2},S_{q},\subseteq,\mathcal{P}(\mathscr{L}))$ such that $W=W_{\mathfrak{S}}$. Since $\mathfrak{S}$ is functionally $2$-valued, $\mathbf{M}\subseteq\{0,1\}^{\mathscr{L}}$.
@@
    We now claim that $W_{\mathfrak{S}}$ is reflexive. To prove this, we first show that $\alpha\in W_{\mathfrak{S}}(\{\alpha\})$ for all $\alpha\in\mathscr{L}$. Suppose the contrary. Then there exists $\alpha\in\mathscr{L}$ such that $\alpha\notin W_{\mathfrak{S}}(\{\alpha\})$. But this would mean that there exists $m\in\mathbf{M}$ such that $m\models_{1}\{\alpha\}$ such that $m\not\models_{2}\{\alpha\}$- a contradiction since $\models_{1}\,=\,\models_{2}$. Thus $\alpha\in W_{\mathfrak{S}}(\{\alpha\})$ for all $\alpha\in\mathscr{L}$. Now, let $\Gamma\subseteq\mathscr{L}$ and $\beta\in\Gamma$, i.e., $\{\beta\}\subseteq\Gamma$. Since $W=W_{\mathfrak{S}}$ and $W$ is monotonic, $W_{\mathfrak{S}}$ is monotonic as well. This implies that $W_{\mathfrak{S}}(\{\beta\})\subseteq W_{\mathfrak{S}}(\Gamma)$. Since $\beta\in W_{\mathfrak{S}}(\{\beta\})$, we must have $\beta\in W_{\mathfrak{S}}(\Gamma)$. So $\beta\in W_{\mathfrak{S}}(\Gamma)$ for all $\beta\in\Gamma$. Consequently, $W_{\mathfrak{S}}$, which implies that $W$ is reflexive as well -- a contradiction. This establishes the theorem. ∎
@@
    Since $W=W_{\mathfrak{S}}$, this shows that $W(W(\Sigma))\subseteq W(\Sigma)$. Now, by reflexivity of $W$, we have $W(\Sigma)\subseteq W(W(\Sigma))$. This shows that $W(W(\Sigma))=W(\Sigma)$ for all $\Sigma\subseteq\mathscr{L}$. Thus, $W$ is idempotent -- a contradiction. This completes the proof. ∎
@@ 3.3 $s_{\kappa}$-consequence Operators
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
    Then, |×$|\mathscr{L}\times\{0\}|=|\mathscr{L}\times\{1\}|=\kappa$. So, there exist a bijection $\sigma:\mathscr{L}\to(\mathscr{L}\times\{0\})\cup(\mathscr{L}\times\{1\})$. Let $X=\sigma^{-1}(\mathscr{L}\times\{0\})$ and $Y=\sigma^{-1}(\mathscr{L}\times\{1\})$. Now,
@@
[…]
@@
[…]
@@
[…]
@@ 3. $\bullet$
[…]
@@
[…]
@@ 3.4 $r_{\kappa}$-consequence Operators
      Since $W$ is not a $q$-consequence operator, by Theorem 3.6(5) there exists $\Gamma\subseteq\mathscr{L}$ such that $W(\Gamma)\not\subseteq\Gamma$. Otherwise, for any $\Gamma\subseteq\mathscr{L}$, $W(\Gamma)\cup\Gamma\subseteq\Gamma$, and hence $W(\Gamma)\cup\Gamma=\Gamma$ which implies that $W(W(\Gamma)\cup\Gamma)=W(\Gamma)$. This, however, would imply that $W$ satisfies quasi-closure, and so is a $q$-consequence operator -- a contradiction.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
    We note that, since $\mathscr{L}$ is infinite, there exists $\beta\in\mathscr{L}$ such that $\alpha\neq\beta$. Now, since |$|\{\beta\}|=1<\kappa$, $W(\{\beta\})=\{\alpha\}$. Then, as $\{\beta\}\not\subseteq W(\{\beta\})=\{\alpha\}$, $W$ is non-reflexive. Since $|\emptyset|=0<\kappa$, $W(\emptyset)=\{\alpha\}\not\subseteq\emptyset$. So, $W$ satisfies non-quasi-closure as well.
[…]
@@
[…]
@@
[…]
@@
    If $\mathfrak{S}$ is functionally $2$-valued, then along the same lines as in the proof of Theorem 3.10, we are done. We skip this, and move on to where $\mathfrak{S}$ is a functionally $3$-valued functional semantics.
@@
[…]

### v2

@@ 3. $\bullet$
    Let $(\mathscr{L},W)$ be a logical structure induced by a non-reflexive $q$-consequence operator. Note that, by Theorem 3.8, it is functionally $3$-valued. Suppose that it is not inferentially $3$-valued. So, there exists some functionally $2$-valued semantics for $\mathscr{L}$, say $\mathfrak{S}=(\textbf{M},\models_{1},\models_{2},S_{q},\subseteq,\mathcal{P}(\mathscr{L}))$ such that $W=W_{\mathfrak{S}}$. Since $\mathfrak{S}$ is functionally $2$-valued, without loss of generality, we may assume that $\mathbf{M}\subseteq\{0,1\}^{\mathscr{L}}$.
@@
    We now claim that $W_{\mathfrak{S}}$ is reflexive. To prove this, we first show that $\alpha\in W_{\mathfrak{S}}(\{\alpha\})$ for all $\alpha\in\mathscr{L}$. Suppose the contrary. Then there exists $\alpha\in\mathscr{L}$ such that $\alpha\notin W_{\mathfrak{S}}(\{\alpha\})$. But this would mean that there exists $m\in\mathbf{M}$ such that $m\models_{1}\{\alpha\}$ such that $m\not\models_{2}\{\alpha\}$ - a contradiction since $\models_{1}\,=\,\models_{2}$. Thus $\alpha\in W_{\mathfrak{S}}(\{\alpha\})$ for all $\alpha\in\mathscr{L}$. Now, let $\Gamma\subseteq\mathscr{L}$ and $\beta\in\Gamma$, i.e., $\{\beta\}\subseteq\Gamma$. Since $W=W_{\mathfrak{S}}$ and $W$ is monotonic, $W_{\mathfrak{S}}$ is monotonic as well. This implies that $W_{\mathfrak{S}}(\{\beta\})\subseteq W_{\mathfrak{S}}(\Gamma)$. Since $\beta\in W_{\mathfrak{S}}(\{\beta\})$, we must have $\beta\in W_{\mathfrak{S}}(\Gamma)$. So $\beta\in W_{\mathfrak{S}}(\Gamma)$ for all $\beta\in\Gamma$. Consequently, $W_{\mathfrak{S}}$, which implies that $W$ is reflexive as well - a contradiction. This establishes the theorem. ∎
@@
    Since $W=W_{\mathfrak{S}}$, this shows that $W(W(\Sigma))\subseteq W(\Sigma)$. Now, by reflexivity of $W$, we have $W(\Sigma)\subseteq W(W(\Sigma))$. This shows that $W(W(\Sigma))=W(\Sigma)$ for all $\Sigma\subseteq\mathscr{L}$. Thus, $W$ is idempotent - a contradiction. This completes the proof. ∎
@@ 3.3 $s_{\kappa}$-consequence Operators
    Definition 3.21 ($\textsc{Anti-reflexivity}^{\mathsf{K}}_{\kappa}$).
    Given a logical structure $(\mathscr{L},W)$, a cardinal $\kappa$ and $\mathsf{K}\subseteq\mathcal{P}(\mathscr{L})$, $W$ is said to be $\textit{anti-reflexive}^{\mathsf{K}}_{\kappa}$ (anti-reflexive relative to $\mathsf{K}$ that is internally $\kappa$) or to satisfy $\textit{anti-reflexivity}^{\mathsf{K}}_{\kappa}$ if for every $\Gamma\in\mathsf{K}$, $\Gamma$ satisfies anti-reflexivity, i.e., $\Gamma\subseteq\mathscr{L}\setminus W(\Gamma)$ for all $\Gamma\in\mathsf{K}$. In this case, $(\mathscr{L},W)$ is also said to $\textit{be anti-reflexive}^{\mathsf{K}}_{\kappa}$.
@@
    Let $\mathscr{L}$ be a set, $W:\mathcal{P}(\mathscr{L})\to\mathcal{P}(\mathscr{L})$ and $\kappa$ a cardinal. Then $W$ is said to be an $s^{\mathsf{K}}_{\kappa}$-consequence operator on $\mathscr{L}$ if it is monotonic and $\text{anti-reflexive}^{\mathsf{K}}_{\kappa}$. A logical structure $(\mathscr{L},W)$ is said to be of $s^{\mathsf{K}}_{\kappa}\textit{-type}$ if $W$ is an $s^{\mathsf{K}}_{\kappa}$-consequence on $\mathscr{L}$.
    $W$ is said to be an $s_{\kappa}$-consequence operator on $\mathscr{L}$ if $W$ is an $s^{\mathsf{K}}_{\kappa}$-consequence operator on $\mathscr{L}$ for some $\mathsf{K}\subseteq\mathscr{L}$ that is internally $\kappa$. A logical structure $(\mathscr{L},W)$ is said to be of $s_{\kappa}\textit{-type}$ if $W$ is an $s_{\kappa}$-consequence operator on $\mathscr{L}$.
@@
    (1) Let $\Gamma\in\mathsf{K}$. Then $\Gamma\subseteq\mathscr{L}\setminus W(\Gamma)$ by $\text{anti-reflexivity}^{\mathsf{K}}_{\kappa}$ of $W$. So, by monotonicity, $W(\Gamma)\subseteq W(\mathscr{L}\setminus W(\Gamma))$.
    Now, suppose that $\mathscr{L}\setminus W(\Gamma)\in\mathsf{K}$ as well. Then, for any $\alpha\notin W(\Gamma)$, $\alpha\in\mathscr{L}\setminus W(\Gamma)$, which implies that $\alpha\notin W(\mathscr{L}\setminus W(\Gamma))$ as $W$ is $\text{anti-reflexive}^{\mathsf{K}}_{\kappa}$. Thus, $W(\mathscr{L}\setminus W(\Gamma))\subseteq W(\Gamma)$.
@@
    (3) Suppose there exists $\Gamma\subseteq\mathscr{L}$ and $\Sigma\in\mathsf{K}$ with $\Gamma\cap\Sigma\neq\emptyset$ such that $W(\Sigma)=\Gamma$. Now, since $W$ is $\text{anti-reflexive}^{\mathsf{K}}_{\kappa}$, we have $\Gamma=W(\Sigma)\subseteq\mathscr{L}\setminus\Sigma$, which implies that $\Gamma\cap\Sigma=\emptyset$ - a contradiction. Hence $W(\Sigma)\neq\Gamma$. ∎
@@
    Hence $(\mathscr{L},W)$ is $\text{anti-reflexive}^{\mathsf{K}}_{\kappa}$.
@@
    $(n,m)\in W(\Gamma)$ $\implies(m,n)\in\Gamma$ $\text{(by the definition of}~W)$
    $\implies(m,n)\in\Sigma$ $\text{(since}~\Gamma\subseteq\Sigma)$
    $\implies(n,m)\in W(\Sigma)$ $\text{(by the definition of}~W)$
@@
    Then, $|\mathscr{L}\times\{0\}|=|\mathscr{L}\times\{1\}|=\kappa$. So, there exist a bijection $\sigma:\mathscr{L}\to(\mathscr{L}\times\{0\})\cup(\mathscr{L}\times\{1\})$. Let $X=\sigma^{-1}(\mathscr{L}\times\{0\})$ and $Y=\sigma^{-1}(\mathscr{L}\times\{1\})$. Now,
@@
    We now claim that for all $\Gamma\in\mathsf{K}$, $\Gamma\subseteq\mathscr{L}\setminus W_{\lambda}(\Gamma)$. Let $\Gamma\in\mathsf{K}$. Then, $|\Gamma|\geq\lambda$ and $\Gamma\subseteq X$ or $\Gamma\subseteq Y$. Without loss of generality, suppose that $\Gamma\subseteq X$. Then, $W_{\lambda}(\Gamma)=f(\Gamma)\subseteq Y$. Now, as $X\cap Y=\emptyset$, $\Gamma\cap W_{\lambda}(\Gamma)=\emptyset$, i.e., $\Gamma\subseteq\mathscr{L}\setminus W_{\lambda}(\Gamma)$. Thus $W_{\kappa}$ is $\text{anti-reflexive}^{\mathsf{K}}_{\lambda}$.
@@
    Let $(\mathscr{L},W)$ be a non-reflexive and monotonic logical structure. Then, by non-reflexivity, there exists $\Gamma\subseteq\mathscr{L}$ such that $\Gamma\not\subseteq W(\Gamma)$. Let $\alpha\in\Gamma$ such that $\alpha\notin W(\Gamma)$. Consequently, since $W$ is monotonic, $\alpha\notin W(\{\alpha\})$ as well. It is now easy to see that $W$ is $\text{anti-reflexive}_{1}^{\{\alpha\}}$. Since $W$ is monotonic as well, $W$ is an $s^{\{\alpha\}}_{1}-$type logical structure, and hence $(\mathscr{L},W)$ is of $s_{1}$-type. ∎
@@
    First, we show that $W_{\mathfrak{S}}$ is $\text{anti-reflexive}^{\mathsf{K}}_{\kappa}$. Suppose the contrary, and let $\Gamma\in\mathsf{K}$ such that $\Gamma\cap W_{\mathfrak{S}}(\Gamma)\neq\emptyset$. Let $\alpha\in\Gamma$ such that $\alpha\in W_{\mathfrak{S}}(\Gamma)$. Now, since $\alpha\in\Gamma$ and $\Gamma\in\mathsf{K}$, by (2), there exists $m\in\mathbf{M}$ such that $m\models_{1}\Gamma$ but $m\not\models_{2}\{\alpha\}$. This, however, implies that $\alpha\notin W_{\mathfrak{S}}(\Gamma)$ - a contradiction. Hence, $W_{\mathfrak{S}}$ is $\text{anti-reflexive}^{\mathsf{K}}_{\kappa}$. Since $\mathfrak{S}$ is strongly granular, $W_{\mathfrak{S}}$ is monotonic, by Theorem 3.1. Thus, $W_{\mathfrak{S}}$ is an $s^{\mathsf{K}}_{\kappa}$-consequence operator. ∎
@@ 3. $\bullet$
    $\implies W(\Gamma)\subseteq W(\Sigma)$ $\text{(since}~W~\text{is monotonic)}$
    $\implies\alpha\in W(\Sigma)$ $\text{(since}~\alpha\in W(\Gamma))$
@@
    Suppose the contrary and let $\Gamma\in\mathsf{K}$ and $\alpha\in\Gamma$ such that $\pi_{1}(\models_{1}\cap\,(\mathbf{M}\times\{\Gamma\}))\subseteq\,\pi_{1}(\models_{2}\cap\,(\mathbf{M}\times\{\{\alpha\}\}))$. Now, since $W$ is $\text{anti-reflexive}^{\mathsf{K}}_{\kappa}$, $\Gamma\in\mathsf{K}$ and $\alpha\in\Gamma$, $\alpha\notin W(\Gamma)$. Then, as $W=W_{\mathfrak{S}}$, $\alpha\notin W_{\mathfrak{S}}(\Gamma)$ as well. So, there exists $m\in\mathbf{M}$ such that $m\models_{1}\Gamma$ but $m\not\models_{2}\{\alpha\}$. Now, $m\not\models_{2}\{\alpha\}$ implies that $m\notin\pi_{1}(\models_{2}\cap\,(\mathbf{M}\times\{\{\alpha\}\}))$, and since $\pi_{1}(\models_{1}\cap\,(\mathbf{M}\times\{\Gamma\}))\subseteq\,\pi_{1}(\models_{2}\cap\,(\mathbf{M}\times\{\{\alpha\}\}))$, it follows that $m\not\models_{1}\Gamma$ - a contradiction. Consequently, $\pi_{1}(\models_{1}\cap\,(\mathbf{M}\times\{\Gamma\}))\not\subseteq\,\pi_{1}(\models_{2}\cap\,(\mathbf{M}\times\{\{\alpha\}\}))$ for all $\Gamma\cup\{\alpha\}\subseteq\mathscr{L}$ with $\Gamma\in\mathsf{K}$ and $\alpha\in\Gamma$. ∎
@@ 3.4 $r_{\kappa}$-consequence Operators
      Since $W$ is not a $q$-consequence operator, by Theorem 3.6(5) there exists $\Gamma\subseteq\mathscr{L}$ such that $W(\Gamma)\not\subseteq\Gamma$. Otherwise, for any $\Gamma\subseteq\mathscr{L}$, $W(\Gamma)\cup\Gamma\subseteq\Gamma$, and hence $W(\Gamma)\cup\Gamma=\Gamma$ which implies that $W(W(\Gamma)\cup\Gamma)=W(\Gamma)$. This, however, would imply that $W$ satisfies quasi-closure, and so is a $q$-consequence operator - a contradiction.
@@
      Since $W$ is monotonic, but not an $s_{\kappa}$-consequence operator, it is not $\text{anti-reflexive}^{\mathsf{K}}_{\kappa}$ for any $\mathsf{K}\subseteq\mathcal{P}(\mathscr{L})$ that is internally $\kappa$. This is the same as saying that for all $\Gamma\subseteq\mathscr{L}$ with $|\Gamma|\geq\kappa$, $\Gamma\not\subseteq\mathscr{L}\setminus W(\Gamma)$.
@@
      For all $\Gamma\subseteq\mathscr{L}$, with $|\Gamma|\geq\kappa$, $\Gamma\not\subseteq\mathscr{L}\setminus W(\Gamma)$. ($\text{Non-anti-reflexivity}_{\kappa}$)
@@
    A logical structure $(\mathscr{L},W)$ is said to be of $r_{\kappa}\textit{-type}$ if $W$ is an $r_{\kappa}$-consequence operator on $\mathscr{L}$.
@@
    Suppose the contrary. Let $(\mathscr{L},W)$ be a logical structure such that $W$ is an $r_{1}$-consequence operator. Since $W$ is non-reflexive, there exists $\Gamma\subseteq\mathscr{L}$ such that $\Gamma\not\subseteq W(\Gamma)$. Therefore, there exists $\alpha\in\Gamma$ such that $\alpha\notin W(\Gamma)$. Since $W$ is monotonic, this implies that $\alpha\notin W(\{\alpha\})$ as well. However, $\alpha\notin W(\{\alpha\})$ implies that $\alpha\in\mathscr{L}\setminus W(\{\alpha\})$, i.e., $\{\alpha\}\subseteq\mathscr{L}\setminus W(\{\alpha\})$, contradicting the fact that $W$ satisfies $\text{non-anti-reflexivity}_{1}$. Thus $(\mathscr{L},W)$ cannot be of $r_{1}$-type. Since $(\mathscr{L},W)$ was chosen arbitrarily, this completes the proof. ∎
@@
    We note that, since $\mathscr{L}$ is infinite, there exists $\beta\in\mathscr{L}$ such that $\alpha\neq\beta$. Now, since $|\{\beta\}|=1<\kappa$, $W(\{\beta\})=\{\alpha\}$. Then, as $\{\beta\}\not\subseteq W(\{\beta\})=\{\alpha\}$, $W$ is non-reflexive. Since $|\emptyset|=0<\kappa$, $W(\emptyset)=\{\alpha\}\not\subseteq\emptyset$. So, $W$ satisfies non-quasi-closure as well.
    Next, let $\Gamma\subseteq\mathscr{L}$ such that $|\Gamma|\geq\kappa$. Then, $W(\Gamma)=\Gamma\cup\{\alpha\}$, and hence $\Gamma\not\subseteq\mathscr{L}\setminus W(\Gamma)$. Thus $W$ satisfies $\text{non-anti-reflexivity}_{\kappa}$.
@@
    It is worth pointing out that the $r_{\kappa}$-consequence operator $W$ in the above example is not an $r_{\lambda}$-consequence operator for any cardinal $\lambda$ such that $2\leq\lambda<\kappa$. To see this, let $\lambda$ be a cardinal such that $2\leq\lambda<\kappa$ and $\Gamma_{\lambda}$ a set of cardinality $\lambda$ such that $\alpha\notin\Gamma_{\lambda}$. The existence of such a $\Gamma_{\lambda}$ can be justified as follows. If $\lambda$ is infinite, then we choose any set $\Gamma$ of cardinality $\lambda$ (which exists by Theorem 3.25), and take $\Gamma\setminus\{\alpha\}$ to be $\Gamma_{\lambda}$. On the other hand, if $\lambda$ is finite, we consider any set $\Gamma$ of cardinality $\lambda$, $\beta\notin\Gamma$ (such a $\beta$ exists since $\mathscr{L}$ is infinite but $\Gamma$ is finite) and define $\Gamma_{\lambda}=(\Gamma\setminus\{\alpha\})\cup\{\beta\}$. Then, as $|\Gamma_{\lambda}|=\lambda<\kappa$, $W(\Gamma_{\lambda})=\{\alpha\}$, which implies that $\Gamma_{\lambda}\subseteq\mathscr{L}\setminus W(\Gamma_{\lambda})=\mathscr{L}\setminus\{\alpha\}$. This shows that $W$ does not satisfy $\text{non-anti-reflexivity}_{\lambda}$. Thus, $W$ is not an $r_{\lambda}$-consequence operator on $\mathscr{L}$ for any cardinal $\lambda$ such that $2\leq\lambda<\kappa$. One might compare this with Example 3.26.
@@
    Finally, let $\Lambda\subseteq\mathscr{L}$ such that $|\Lambda|\geq\kappa$. Then, by (3), there exists $\lambda\in\Lambda$ such that $\pi_{1}(\models_{1}\cap\,(\mathbf{M}\times\{\Lambda\}))\,\subseteq\,\pi_{1}(\models_{2}\cap\,(\mathbf{M}\times\{\{\lambda\}\}))$. We choose any such $\lambda$ and $m\in\mathbf{M}$ such that $m\models_{1}\Lambda$. Then, since $\pi_{1}(\models_{1}\cap\,(\mathbf{M}\times\{\Lambda\}))\,\subseteq\,\pi_{1}(\models_{2}\cap\,(\mathbf{M}\times\{\{\lambda\}\}))$, $m\models_{2}\{\lambda\}$. Thus, $\lambda\in\Lambda\cap W_{\mathfrak{S}}(\Lambda)$ and hence $\Lambda\not\subseteq\mathscr{L}\setminus W_{\mathfrak{S}}(\Lambda)$. Therefore, $W_{\mathfrak{S}}$ satisfies $\text{non-anti-reflexivity}_{\kappa}$ as well.
@@
    If $\mathfrak{S}$ is functionally $2$-valued, then along the same lines as in the proof of Theorem 3.10, we are done. We skip this, and move on to where $\mathfrak{S}$ is a functionally $3$-valued functional semantics. In this case, without loss of generality, we may assume that $\mathbf{M}\subseteq\{0,1,2\}^{\mathscr{L}}$.
@@
    Either $D_{1}\cap D_{2}=\emptyset$ or $D_{1}\cap D_{2}\neq\emptyset$. Suppose $D_{1}\cap D_{2}=\emptyset$. Let $\Gamma\subseteq\mathscr{L}$ such that $|\Gamma|\geq\kappa$ and $\alpha\in\Gamma$. Moreover, let $m\in\mathbf{M}$ such that $m\models_{1}\Gamma$, i.e., $m(\Gamma)\subseteq D_{1}$. Then, as $D_{1}\cap D_{2}=\emptyset$, $m(\alpha)\notin D_{2}$, i.e., $m\not\models_{2}\{\alpha\}$. In other words, $\alpha\notin W_{\mathfrak{S}}(\Gamma)$. Thus $\Gamma\subseteq\mathscr{L}\setminus W_{\mathfrak{S}}(\Gamma)$ for all $|\Gamma|\geq\kappa$, which implies that $W$ is $\text{anti-reflexive}^{\mathsf{K}}_{\kappa}$. This is a contradiction to the assumption that $W$ is a $r_{\kappa}$-consequence operator. So, $D_{1}\cap D_{2}\neq\emptyset$.